# Assessing (im)balance in signed brain networks

### v1

@@ Abstract
Many complex systems - be they financial, natural or social - are composed by units - such as stocks, neurons or agents - whose joint activity can be represented as a multivariate time series. An issue of both practical and theoretical importance concerns the possibility of inferring the presence of a static relationships between any two units solely from their dynamic state. The present contribution aims at providing an answer within the frame of traditional hypothesis testing. Briefly speaking, our suggestion is that of linking any two units if behaving in a sufficiently similar way. To achieve such a goal, we project a multivariate time series onto a signed graph, by i) comparing the empirical properties of the former with those expected under a suitable benchmark and ii) linking any two units with a positive (negative) edge in case the corresponding series share a significantly large number of concordant (discordant) values. To define our benchmarks, we adopt an information-theoretic approach that is rooted into the constrained maximisation of Shannon entropy, a procedure inducing an ensemble of multivariate time series that preserves some of the empirical properties on average while randomising everything else. We showcase the possible applications of our method by addressing one of the most timely issues in the domain of neurosciences, i.e. that of determining if brain networks are frustrated or not - and, in case, to what extent. As our results suggest, this is indeed the case, the structure of the negative subgraph being more prone to inter-subject variability than the complementary, positive subgraph. At the mesoscopic level, instead, the minimisation of the Bayesian Information Criterion instantiated with the Signed Stochastic Block Model reveals that brain areas gather into modules aligning with the statistical variant of the Relaxed Balance Theory.
@@ I Introduction
Time series are ubiquitous and versatile representations of dynamical complex systems, be they financial, natural or social. A major challenge is that of inferring the underlying dependencies between units from the (often non-linear) dynamics characterising them. Being capable of transforming a time series into a symbolic sequence, while preserving essential information about the ‘original’ generative process, is of practical relevance, as the dynamical properties of the former can be revealed by focusing on the structural properties of the latter Onnela et al. (2004); Sinha et al. (2010); Bouchaud and Potters (2003). To this aim, sequential data are often represented as traditional graphs, with nodes replacing individual units (such as stocks, neurons or agents) and edges reflecting their activity-driven interactions.
This choice plays a major role in the neuroscientific domain, where network-based approaches have revolutionised our understanding of the brain functioning, especially through the study of functional connectivity networks derived from Blood Oxygen Level Dependent (BOLD) fMRI signals. Rooted in graph theory and statistical physics, this perspective, often referred to as network neuroscience, aims at revealing the organisational principles of cognition and behaviour Kulkarni and Bassett (2025) by modelling the brain as an integrated system of interacting regions Bassett and Sporns (2017).
A review of the approaches that have been proposed so far to extract relationships between nodes directly from the time series of the original units is provided below.
@@ I.1 From time series to proximity and from proximity to graphs
Given a set of time series, the most common approach is that of extracting a proximity matrix from it and, then, a graph from the proximity matrix. Although several definitions of proximity have been proposed [^1] We redirect the interested reader to the reviews Cliff et al. (2023); Liu et al. (2024), where hundreds of proximity measures (such as the dynamic time warping, the mutual information and the transfer entropy) are discussed and compared., the commonest choice is identifying it with the notion of pairwise correlation, as returned by the calculation of the Pearson correlation coefficient
@@
A less popular, but still worth to be mentioned, definition of proximity is the one resting upon the notion of partial correlation Hampson et al. (2002); Fransson and Marrelec (2008). If we indicate the matrix of pairwise correlations with $\mathbf{C}$, the generic entry of the matrix of partial correlations, then, reads $\hat{J}_{ik}=J_{ik}/\sqrt{J_{ii}\cdot J_{kk}}$, with $\mathbf{J}=\mathbf{C}^{-1}$ being the so-called precision matrix. Estimations of the functional connectivity (FC) based on partial correlations are affected by a smaller inter-subject variability Brier et al. (2015); Rahim et al. (2019), as they have been shown to resemble more closely the so-called structural connectivity (SC) patterns Huang et al. (2010); Salvador et al. (2005); Lee et al. (2011); Santucci et al. (2025).
@@ I.1.1 Fixed thresholding algorithms
The simplest approach of this kind is based on the interpretation of a correlation matrix (be it partial or not) as a fully connected weighted matrix. In order to obtain non-trivial structures within such a context, sparsification procedures become, thus, necessary, the simplest one being that of thresholding correlations Bullmore and Sporns (2009). This procedure ultimately boils down to remove the weaker connections - usually the most affected by experimental noise Rubinov and Sporns (2010) - while retaining the stronger ones.
@@
In the neuroscientific literature, one of the earliest attempts of this kind was the one pursued in Stam (2004), where measures of synchronisation between brain regions, derived from magnetoencephalographic (MEG) recordings, were considered to extract frequency band-specific networks by fixing their average degree; a similar attempt was pursued in Micheloyannis et al. (2009), where electroencephalographic (EEG) recordings were, instead, considered.
@@
In the financial literature, the network resulting from the application of a threshold is also known with the name of asset graph Sinha et al. (2010); Onnela et al. (2004); Heimo et al. (2009). In Onnela et al. (2004), the authors derived an asset graph from the correlation matrix of a set of stocks by filling an empty graph, edge by edge, starting from the one corresponding to the largest correlation and stopping at the $N$-th largest correlation (with $N$ being the number of nodes). Upon doing so, the considered asset graph is comparable in size with the most popular asset tree, although being less constrained by requests such as the one of forbidding cycles.
As a last comment, we report the concern expressed by the authors of MacMahon and Garlaschelli (2015), according to whom thresholding may prevent the detection of a modular organisation, as the correlations within modules may be cut out even if significantly stronger than those between modules.
@@
Since letting the network density drive the thresholding procedure appears as arbitrary as setting a threshold directly on correlations, it has been proposed to determine it by requiring that the resulting network remains connected. Such a data-driven procedure removes the weaker correlations until the percolation threshold is reached, i.e. the value in correspondence of which the resulting network disconnects Alexander-Bloch et al. (2010); Gallos et al. (2012). Although this kind of analysis has been popularised with the name of percolation analysis, the aforementioned value is, technically speaking, the connectivity threshold.
@@
A different, yet related, criterion prescribes to adopt the threshold that maximises the value $\text{APL}_{\text{ER}}=\ln N/\ln\overline{k}$, with $\overline{k}=\sum_{i=1}^{N}k_{i}/N$, i.e. the (expected) average path length of an Erdös-Rényi network whose connectance matches the empirical one Hayasaka and Laurienti (2010); Marangolo et al. (2016); alternatively, one may adopt the threshold [^2] As the authors themselves explicitly state, their criteria individuate a whole range of link densities, more precisely from 0.4$\simeq$ 0.4≃ 0.4 to 0.5$\simeq$ 0.5≃ 0.5: below 0.4$\simeq$ 0.4≃ 0.4 graphs fragment and above 0.5$\simeq$ 0.5≃ 0.5 their topology becomes increasingly random. induced by the request of having a small-world network with no isolated nodes Bassett et al. (2008); Lynall et al. (2010).
@@
If the previous approaches prescribe to stop pruning correlations once the percolation threshold is reached, a second approach focuses on the mesoscopic organisation of the network and prescribes to consider a correlation-induced distance333 Examples are provided the one reading $d_{ij}=\sqrt{2(1-C_{ij})}$, mapping distances between time series into the interval [0,2][0,2][ 0 , 2 ], and the one reading $d_{ij}=\sqrt{(1-C_{ij})/2}$, mapping distances between time series into the interval [0,1][0,1][ 0 , 1 ]. In the first case, 0 corresponds to linearly correlated time series, 2 corresponds to anti-correlated time series and $\sqrt{2}$ corresponds to the case where a linear correlation is not found; in the second case, 0 corresponds to linearly correlated time series, 1 corresponds to anti-correlated time series and $\sqrt{2}/2$ corresponds to the case where a linear correlation is not found., draw the corresponding dendrogram [^4] A dendrogram is constructed by clustering objects in a hierarchical fashion: to this aim, a distance between clusters is needed. Some of these ‘linkages’ algorithms induce an ultra-metric distance on the data Rammal et al. (1986) that has been argued to induce distances progressively less similar to the original ones as higher levels of the taxonomic tree are climbed: according to this intuition, such a method is more reliable to determine the low-level structure of the tree (induced by the strongest correlations) than the high-level one (induced by the weakest correlations) MacMahon and Garlaschelli (2015); Milligan (1979). and cut it in correspondence of a properly-defined threshold Mantegna and Stanley (1999); Tola et al. (2008). A possible criterion to individuate it is that of maximising the modular character of the resulting network structure, analysed upon running one of the many algorithms that have been proposed to this aim (Louvain, Infomap, Surprise, etc.) Newman (2006); Bordier et al. (2017); Marchese et al. (2022).
@@
Examples from the financial literature are provided by Di Matteo et al. (2004), whose authors cut the dendrogram induced by 343434 weekly time series of interest rates at a distance which is halfway between the one induced by $C_{ij}=0$ and the one induced by $C_{ij}=1$, thus recovering 666 clusters throughout the whole period considered there, i.e. 198219821982-199719971997, and Tola et al. (2008), whose authors compare the performance of a similar clustering algorithm with that of an algorithm based on Random Matrix Theory in singling out groups of assets with comparable returns.
@@ I.1.2 Progressive thresholding algorithms
Since individuating a unique threshold rises doubts about the arbitrariness of such a choice - which may significantly impact the results - alternatives have been proposed to make this approach a bit more principled, such as monitoring the changes in the resulting network structure upon rising the threshold itself. The earliest attempt, pursued in Eguíluz et al. (2005), consisted in identifying three threshold values, i.e. 0.60.60.6, 0.70.70.7 and 0.80.80.8: topological traits (averaged over the three sets of threshold-specific configurations) such as the scale-freeness of the distribution of functional connections and the small-worldness of the resulting network structure were, however, found to be robust against progressive thresholding. In Honey et al. (2009), the authors investigated how the relationships between FC and SC evolve upon progressively removing all FC values, the aim being that of quantifying the extent to which the SC values can be inferred from the ‘surviving’ FC ones. Studying the progressive engagement of brain regions has been proven useful to reveal the hierarchical character of the mouse brain Bardella et al. (2016) as well as distinguish healthy subjects from schizophrenic patients Mastrandrea et al. (2021).
@@ I.1.3 Filtering algorithms
Filtering correlations via the MST. The most popular filtering approach returns the so-called Maximum Spanning Tree (MST), a technique retaining the $N-1$ largest correlations and ensuring that i) loops are discarded; ii) each node is reachable from any other node via a path that is the widest between its endpoints; iii) the total weight is greater than, or equal to, the total weight of any other spanning tree. It can be obtained upon running the Kruskal’s algorithm and is often employed to distinguish the connector areas (i.e. the most central ones) from the provincial areas (i.e. the most peripheral ones) Rammal et al. (1986); Mantegna and Stanley (1999); Caldarelli et al. (2012); Mastrandrea et al. (2017, 2021).
@@
Both the MST and the MSF are rather different from their random counterparts Bardella et al. (2016), their structural organisation providing a topological signature of the arising of diseases such as schizophrenia Mastrandrea et al. (2021). Coming to the financial literature, the authors of Bonanno et al. (2004) filter the correlation matrix induced by a set of daily time series of stock returns from the NYSE throughout the whole period considered there, i.e. 198719871987-199819981998, by calculating the Minimum Spanning Tree defined by the $N-1$ smallest correlation-induced distances (see footnote 3).
Filtering correlations via the PMFG. An alternative approach, discarding less information, returns the so-called Planar Maximally Filtered Graph (PMFG) Aste et al. (2005); Tumminello et al. (2005). The criterion informing such a method prescribes to retain the correlations ensuring that the resulting network is a planar graph, i.e. a graph that can be drawn with no links crossing each other (i.e. intersecting each other not in correspondence of a node). Since the PMFG always contains the MST, the former complements the latter by providing additional (and not just different) information. The PMFG has also been described as the simplest instance of a more general procedure embedding high-dimensional data into lower-dimensional manifolds having a controllable number of ‘holes’ called ‘genus’. The PMFG corresponds to the case in which the genus is zero Tumminello et al. (2005).
Although versatile - the method has been extended in a variety of ways, e.g. to produce a nested hierarchy of time series Song et al. (2011, 2012); Morales et al. (2014) - it suffers from some degree of arbitrariness too, affecting the properties of the postulated, approximating structure: the value of the genus must, in fact, be fixed a priori; besides, there is no obvious reason why time series should find a natural embedding into a bidimensional plane.
@@ I.1.4 Testing correlations
A more principled approach to extract information from a correlation matrix is that of comparing (some of) its properties with the ones expected under a properly defined benchmark, the aim of this class of techniques being that of identifying non-random collective information about the set of time series induced by $\mathbf{C}$. Quite interestingly, these techniques do not discard negative correlations, known to be a consequence of spatial and temporal heterogeneities induced by non-trivial, neural-mediated, hemodynamic mechanisms and contributing to the stability of resting-state brain networks, by affecting the synchronisation of neural activity Saberi et al. (2021).
Although isolated attempts exist - such as the one proposed in Salvador et al. (2005), whose authors test the null hypothesis that the average value of the distribution of partial correlations, for each pair of areas and across subjects, is zero - methods of this kind can be split into those based on Random Matrix Theory (RMT) and those based on Entropy Maximization (EM).
Null models based on RMT. One of the most popular models studied in the context of RMT is the Wishart ensemble Wishart (1948); Livan et al. (2018) which, in its simplest form, leads to the Marcenko-Pastur distribution Marčenko and Pastur (1967), i.e. the distribution of the eigenvalues of a correlation matrix induced by $N$ completely random time series of duration $T$ in the limits $N\rightarrow+\infty$ and $T\rightarrow+\infty$ with $1<T/N<+\infty$, and reads
@@
with $\sqrt{\lambda_{\pm}}=1\pm\sqrt{N/T}$ and $\rho(\lambda)=0$ otherwise. As a consequence, while the eigenvalues of an empirical correlation matrix falling within the range $[\lambda_{-},\lambda_{+}]$ can be imputable to random noise, any other eigenvalue can be deemed as representing meaningful structure in the data. As a consequence, the random component of any empirical correlation matrix can be identified with the tensor reading
$|\mathbf{C}^{(r)}=\sum_{i:\lambda_{i}\leq\lambda_{+}}\lambda_{i}\ket{v_{i}}\bra{v_{i}}$ | (4)
where |$\ket{v_{i}}$ is the eigenvector associated with the eigenvalue [^7] The eigenvalues smaller than $\lambda_{-}$ are usually also included. $\lambda_{i}$.
A typical feature of the spectrum of empirical correlation matrices is that of having an eigenvalue $\lambda_{m}$ that is (much) larger than any other eigenvalue and whose corresponding eigenvector |$\ket{v_{m}}$ has all positive components: such a ‘market mode’ contributes to the random component of an empirical correlation matrix $\mathbf{C}$ with a tensor reading $|\mathbf{C}^{(m)}=\lambda_{m}\ket{v_{m}}\bra{v_{m}}$ |.
@@
$\mathbf{C}^{(g)}=\mathbf{C}-\mathbf{C}^{(r)}-\mathbf{C}^{(m)}=\sum_{i:\lambda_{+}<\lambda_{i}<\lambda_{m}}\lambda_{i}\ket{v_{i}}\bra{v_{i}},$ (5)
@@
with $|\mathbf{C}|=\sum_{i=1}^{N}\sum_{j=1}^{N}C_{ij}$ and whose maximisation can be proven to lead to clusters of nodes that are positively correlated while nodes belonging to different clusters are negatively correlated MacMahon and Garlaschelli (2015).
In Plerou et al. (2002), the authors employ models of this kind to analyse correlations of time series of stock market returns: by comparing them with null models based on RMT, the authors find that most eigenvalues are compatible with statistical noise, those capturing meaningful correlations (and being linked to economic factors such as market-wide movements or sectoral influences) being the few, large, deviating ones. In Utsugi et al. (2004), the authors find that the distribution of the eigenvalues of the correlation matrix induced by a set of daily time series of stock prices from the TSE is stable, throughout the whole period of time considered there, i.e. 199319931993-200120012001. In Potters et al. (2005), the authors explore how RMT can be applied to better estimate financial risk. In Zema et al. (2025), the authors build upon such a scheme by developing a novel approach for portfolio management.
Null models based on EM. A different, yet related, approach considers null models induced by entropy maximisation Masuda et al. (2018): in this case, one seeks to maximise the differential entropy
[…]
@@
In Kojaku and Masuda (2019), the authors employ the same null model to filter covariance matrices, inducing networks in which any two nodes are linked if their correlation is found to be significantly larger than expected: to this aim, the matrix $\bm{\Delta}=\mathbf{C}-\mathbf{C}^{(r)}$, defined as the difference between the empirical and the random covariance matrix, is numerically determined by maximising the log-likelihood function [^8] To avoid overfitting, the log-likelihood is penalised by an adaptive-Lasso term whose parameters are determined by minimising the Bayesian Information Criterion (BIC). A package named ‘SCOLA’ (an acronym stating for ‘Sparse Networks Construction from Correlational Data with Lasso’) has been also released. $\mathcal{L}=\ln P(\mathbf{X}|\mathbf{C})$.
Bayesian inference. Another way for inferring a sparse graph from a set of time series, while avoiding thresholding, is that of inferring a precision matrix with $\ell_{1}$ penalization, where the value of the lasso regularisation parameter is usually set by cross-validation. This is equivalent to choosing the precision matrix that maximises the posterior distribution given the data, using a Gaussian likelihood and a prior encoding the $\ell_{1}$ constraint Ibáñez Berganza et al. (2023); Santucci et al. (2025); Varoquaux and Craddock (2013). Such an approach leads to a sparse, positive-definite precision matrix, interpretable as an adjacency matrix.

### v3

@@ Abstract
Many complex systems - be they financial, natural, or social - are composed of units - such as stocks, neurons, or agents - whose joint activity can be represented as a multivariate time series. An issue of both practical and theoretical importance concerns the possibility of inferring the presence of a static relationship between any two units solely from their dynamic state. The present contribution aims at tackling such an issue within the frame of traditional hypothesis testing: briefly speaking, our suggestion is that of linking any two units if behaving in a sufficiently similar way. To achieve such a goal, we project a multivariate time series onto a signed graph by i) comparing the empirical properties of the former with those expected under a suitable benchmark and ii) linking any two units with a positive (negative) edge in case the corresponding series shares a significantly large number of concordant (discordant) values. To define our benchmarks, we adopt an information-theoretic approach that is rooted into the constrained maximisation of Shannon entropy, a procedure inducing an ensemble of multivariate time series that preserves some of the empirical properties on average, while randomising everything else. We showcase the possible applications of our method by addressing one of the most timely issues in the domain of neurosciences, i.e. that of determining if brain networks are frustrated or not, and, if so, to what extent. As our results suggest, this is indeed the case, with the major contribution to the underlying negative subgraph coming from the subcortical structures (and, to a lesser extent, from the limbic regions). At the mesoscopic level, the minimisation of the Bayesian Information Criterion, instantiated with the Signed Stochastic Block Model, reveals that brain areas gather into modules aligning with the statistical variant of the Relaxed Balance Theory.
@@ I Introduction
Time series are ubiquitous and versatile representations of dynamical complex systems, be they financial, natural, or social. A major challenge is that of inferring the underlying dependencies between units from the (often non-linear) dynamics characterising them. Being capable of transforming a time series into a symbolic sequence, while preserving essential information about the ‘original’ generative process, is of practical relevance, as the dynamical properties of the former can be revealed by focusing on the structural properties of the latter Onnela2004; Sinha2010; Bouchaud2003. To this aim, sequential data are often represented as traditional graphs, with nodes replacing individual units (such as stocks, neurons, or agents) and edges reflecting their activity-driven interactions.
This choice plays a major role in the neuroscientific domain, where network-based approaches have revolutionised our understanding of brain functioning, especially through the study of functional connectivity networks derived from Blood Oxygen Level-Dependent (BOLD) fMRI signals. Rooted in graph theory and statistical physics, this perspective, often referred to as network neuroscience, aims at revealing the organisational principles of cognition and behaviour kulkarni2024principlesbrainnetworkorganization by modelling the brain as an integrated system of interacting regions bassettnetwork2017. In the following subsection, a review of the approaches proposed so far to extract relationships between nodes directly from the time series of the original units is provided.
@@ I.1 From time series to proximity and from proximity to graphs
Given a set of time series, the most common approach is that of extracting a proximity matrix that is subsequently used for defining a graph. Although several definitions of proximity have been proposed [^1] We redirect the interested reader to the reviews Cliff2023; Liu2024, where hundreds of proximity measures (such as the dynamic time warping, the mutual information and the transfer entropy) are discussed and compared., the commonest choice is identifying it with the notion of pairwise correlation, as returned by the calculation of the Pearson correlation coefficient
@@
A less popular, but still worth mentioning, definition of proximity is based on the notion of partial correlation Hampson2002; Fransson2008. If we keep indicating the matrix of pairwise correlations with $\mathbf{C}$, the generic entry of the matrix of partial correlations, then, reads $\hat{J}_{ik}=J_{ik}/\sqrt{J_{ii}\cdot J_{kk}}$, with $\mathbf{J}=\mathbf{C}^{-1}$ being the so-called precision matrix. Estimations of the functional connectivity (FC) based on partial correlations are affected by a smaller inter-subject variability Brier2015; Rahim2019, as they have been shown to resemble more closely the so-called structural connectivity (SC) patterns Huang2010; Salvador2005; Lee2011; Santucci2025.
@@ I.1.1 Fixed thresholding algorithms
The simplest approach of this kind is based on interpreting a correlation matrix (be it partial or not) as a fully-connected, weighted matrix. Non-trivial structures can thus be obtained by implementing sparsification procedures, the simplest of which is thresholding correlations Bullmore2009. This procedure ultimately boils down to remove the weaker connections - usually the most affected by experimental noise Rubinov2010 - while retaining the stronger ones; still, as reported by the authors of MacMahon2015, thresholding may prevent the detection of a modular organisation, as the correlations within modules may be cut out even if significantly stronger than those between modules.
@@
In the neuroscientific literature, one of the earliest algorithms of this kind was the one pursued in Stam2004, where measures of synchronisation between brain regions, derived from magnetoencephalographic (MEG) recordings, were considered and frequency band-specific networks extracted by fixing the average degree; a similar attempt was pursued in Micheloyannis2007, where electroencephalographic (EEG) recordings were, instead, considered.
@@
In the financial literature, the network resulting from the application of a threshold is also known as the asset graph Sinha2010; Onnela2004; Heimo2009. In Onnela2004, the authors derived an asset graph from the correlation matrix of a set of stocks by filling an empty graph, edge by edge, starting from the one corresponding to the largest correlation and stopping at the $N$-th largest correlation (with $N$ being the number of nodes). Upon doing so, the considered asset graph is comparable in size with the most popular asset tree, although being less constrained by requests such as the one of forbidding cycles.
@@
Since letting the network density drive the thresholding procedure appears as arbitrary as setting a threshold directly on correlations, the proposal to determine the resulting network by requiring it to remain connected has been advanced. Such a data-driven procedure removes the weaker correlations until the value in correspondence of which the resulting network disconnects is reached Alexander-Bloch2010; Gallos2012. Although this kind of analysis has been popularised with the name of percolation analysis, the aforementioned value is, technically speaking, the connectivity threshold.
@@
A different, yet related, criterion prescribes to adopt the threshold that maximises the value $\text{APL}_{\text{ER}}=\ln N/\ln\overline{k}$, with $\overline{k}=\sum_{i=1}^{N}k_{i}/N$, i.e. the (expected) average path length of an Erdös-Rényi network whose connectance matches the empirical one Hayasaka2010; Marangolo2016; alternatively, one may adopt the threshold [^2] As the authors themselves explicitly state, their criterion individuate a whole range of link densities, more precisely from $\simeq 0.4$ to $\simeq 0.5$: below $\simeq 0.4$ graphs fragment and above $\simeq 0.5$ their topology becomes increasingly random. induced by the request of having a small-world network with no isolated nodes Bassett2008; Lynall2010.
@@
If the previous approaches prescribe to stop pruning correlations once the connectivity threshold is reached, a second approach focuses on the mesoscopic organisation of the network and prescribes to consider a correlation-induced distance [^3] Examples are provided the one reading $d_{ij}=\sqrt{2(1-C_{ij})}$, mapping distances between time series into the interval $[0,2]$, and the one reading $d_{ij}=\sqrt{(1-C_{ij})/2}$, mapping distances between time series into the interval $[0,1]$. In the first case, 0 corresponds to linearly correlated time series, 2 corresponds to anti-correlated time series and $\sqrt{2}$ corresponds to the case where a linear correlation is not found; in the second case, 0 corresponds to linearly correlated time series, 1 corresponds to anti-correlated time series and $\sqrt{2}/2$ corresponds to the case where a linear correlation is not found., draw the corresponding dendrogram [^4] A dendrogram is constructed by clustering objects in a hierarchical fashion: to this aim, a distance between clusters is needed. Some of these ‘linkages’ algorithms induce an ultra-metric distance on the data Rammal1986 that has been argued to induce distances progressively less similar to the original ones as higher levels of the taxonomic tree are climbed. According to this intuition, such a method is more reliable to determine the low-level structure of the tree (induced by the strongest correlations) than the high-level one (induced by the weakest correlations) MacMahon2015; Milligan1979. and cut it in correspondence of a properly-defined threshold Mantegna1999; Tola2008. A possible criterion to individuate it is that of maximising the modular character of the resulting network structure, analysed upon running one of the many algorithms that have been proposed to this aim (Louvain, Infomap, Surprise, etc.) Newman2006; Bifone2017; Marchese2022.
@@
Examples from the financial literature are provided by DiMatteo2004, whose authors cut the dendrogram induced by $34$ weekly time series of interest rates at a distance which is halfway between the one induced by $C_{ij}=0$ and the one induced by $C_{ij}=1$, thus recovering $6$ clusters throughout the whole period considered there, i.e. 19821982-$1997$, and Tola2008, whose authors compare the performance of a similar clustering algorithm with that of an algorithm based on Random Matrix Theory in singling out groups of assets with comparable returns.
@@ I.1.2 Progressive thresholding algorithms
Since individuating a unique threshold raises doubts about the arbitrariness of such a choice - which may significantly impact the results - alternatives have been proposed to make this approach a bit more principled, such as monitoring the changes in the resulting network structure upon raising the threshold itself. The earliest attempt, pursued in Eguiluz2005, consisted in identifying three threshold values, i.e. $0.6$, $0.7$, and $0.8$: topological traits (averaged over the three sets of threshold-specific configurations) such as the scale-freeness of the distribution of functional connections and the small-worldness of the resulting network structure were, however, found to be robust against progressive thresholding. In Honey2009, the authors investigated how the relationships between FC and SC evolve upon progressively removing all FC values, the aim being that of quantifying the extent to which the SC values can be inferred from the ‘surviving’ FC ones. Studying the progressive engagement of brain regions has been proven useful to reveal the hierarchical character of the mouse brain Bardella2016, as well as to distinguish healthy subjects from schizophrenic patients Mastrandrea2021.
@@ I.1.3 Filtering algorithms
Filtering correlations via the MST. The most popular filtering approach returns the so-called Maximum Spanning Tree (MST), a technique retaining the N1N-1 largest correlations while ensuring that i) loops are discarded; ii) each node is reachable from any other node via a path that is the widest between its endpoints; iii) the total weight is greater than, or equal to, the total weight of any other spanning tree. This method can be applied upon running the Kruskal’s algorithm and is often employed to distinguish the connector areas (i.e. the most central ones) from the provincial areas (i.e. the most peripheral ones) Rammal1986; Mantegna1999; Caldarelli2012; Mastrandrea2017; Mastrandrea2021.
@@
Both the MST and the MSF are rather different from their random counterparts Bardella2016, their structural organisation providing a topological signature of the emergence of diseases such as schizophrenia Mastrandrea2021. Coming to the financial literature, the authors of Bonanno2004 filter the correlation matrix induced by a set of daily time series of stock returns from the NYSE throughout the whole period considered there, i.e. 19871987-$1998$, by calculating the Minimum Spanning Tree defined by the N1N-1 smallest correlation-induced distances (see footnote 3).
Filtering correlations via the PMFG. An alternative approach, discarding less information, returns the so-called Planar Maximally Filtered Graph (PMFG) Aste2005; Tumminello2005. The criterion informing such a method prescribes to retain the correlations ensuring that the resulting network is a planar graph, i.e. a graph that can be drawn with no links crossing each other - more precisely, one in which edges may intersect only at common endpoints (nodes). Since the PMFG always contains the MST, the former complements the latter by providing additional (and not just different) information. The PMFG has also been described as the simplest instance of a more general procedure embedding high-dimensional data into lower-dimensional manifolds with a controllable number of ‘holes’ called ‘genus’. The PMFG corresponds to the case in which the genus is zero Tumminello2005.
Although versatile - the method has been extended in a variety of ways, e.g. to produce a nested hierarchy of time series Song2011; Song2012; Morales2014 - it suffers from some degree of arbitrariness too, affecting the properties of the postulated, approximating structure: the value of the genus must, in fact, be fixed a priori; besides, there is no obvious reason why time series should find a natural embedding into a bidimensional plane.
@@ I.1.4 Testing correlations
A more principled approach to extract information from a correlation matrix is that of comparing (some of) its properties with the ones expected under a properly defined benchmark, the aim of this class of techniques being that of identifying non-random collective information about the set of time series induced by $\mathbf{C}$. Interestingly, these techniques do not discard negative correlations, which are known to be a consequence of spatial and temporal heterogeneities induced by non-trivial, neural-mediated hemodynamic mechanisms affecting the synchronisation of neural activity, thus contributing to the stability of resting-state brain networks Saberi2021.
Although isolated attempts exist - such as the one proposed in Salvador2005, whose authors test the null hypothesis that the average value of the distribution of partial correlations, for each pair of areas and across subjects, is zero - methods of this kind can be split into those based on Random Matrix Theory (RMT) and those based on Entropy Maximization (EM).
Null models based on RMT. One of the most popular models studied in the context of RMT is the Wishart ensemble Wishart1948; Livan2018 which, in its simplest form, leads to the Marcenko-Pastur distribution MP, i.e. the distribution of the eigenvalues of a correlation matrix induced by $N$ completely random time series of duration $T$ in the limits $N\rightarrow+\infty$ and $T\rightarrow+\infty$ with $1<T/N<+\infty$, and reads
@@
with $\sqrt{\lambda_{\pm}}=1\pm\sqrt{N/T}$ and $\rho(\lambda)=0$ otherwise. As a consequence, while the eigenvalues of an empirical correlation matrix falling within the range $\lambda_{-},\lambda_{+}]$ can be imputable to random noise, any other eigenvalue can be deemed as representing meaningful structure in the data. Thus, the random component of any empirical correlation matrix can be identified with the tensor reading
$|\mathbf{C}^{(r)}=\sum_{i:\lambda_{i}\leq\lambda_{+}}\lambda_{i}\ket{v_{i}}\bra{v_{i}}$ (4)
[…]
@@
$|,\mathbf{C}^{(g)}=\mathbf{C}-\mathbf{C}^{(r)}-\mathbf{C}^{(m)}=\sum_{i:\lambda_{+}<\lambda_{i}<\lambda_{m}}\lambda_{i}\ket{v_{i}}\bra{v_{i}}$, (5)
@@
with $|\mathbf{C}|=\sum_{i=1}^{N}\sum_{j=1}^{N}C_{ij}$ and whose maximisation can be proven to lead to clusters of nodes that are positively correlated while nodes belonging to different clusters are negatively correlated MacMahon2015.
In Plerou2002, the authors employ models of this kind to analyse correlations of time series of stock market returns: by comparing them with null models based on RMT, the authors find that most eigenvalues are compatible with statistical noise, those capturing meaningful correlations (and being linked to economic factors such as market-wide movements or sectoral influences) being the few, large, deviating ones. In Utsugi2004, the authors find that the distribution of the eigenvalues of the correlation matrix induced by a set of daily time series of stock prices from the TSE is stable throughout the whole period of time considered there, i.e. 19931993-$2001$. In Potters2005, the authors explore how RMT can be applied to better estimate financial risk. In Zema2024, the authors build upon such a scheme by developing a novel approach for portfolio management.
Null models based on EM. A different, yet related, approach considers null models induced by entropy maximisation Masuda2018: in this case, one seeks to maximise the differential entropy
$S=-\int P(\mathbf{X})\ln P(\mathbf{X})d\mathbf{X},$ (7)
@@
In Kojaku2019, the authors employ the same null model to filter covariance matrices, inducing networks in which any two nodes are linked if their correlation is found to be significantly larger than expected: to this aim, the matrix $\bm{\Delta}=\mathbf{C}-\mathbf{C}^{(r)}$, defined as the difference between the empirical and the random covariance matrix, is numerically determined by maximising the log-likelihood function [^8] To avoid overfitting, the log-likelihood is penalised by an adaptive-Lasso term whose parameters are determined by minimising the Bayesian Information Criterion (BIC). A package named ‘SCOLA’ (an acronym standing for ‘Sparse Networks Construction from Correlational Data with Lasso’) has also been released. $\mathcal{L}=\ln P(\mathbf{X}|\mathbf{C})$.
Bayesian inference. A third way for inferring a sparse graph from a set of time series, while avoiding thresholding, is that of inferring a precision matrix with $\ell_{1}$ penalisation, where the value of the lasso regularisation parameter is usually set by cross-validation. This is equivalent to choosing the precision matrix that maximises the posterior distribution given the data, using a Gaussian likelihood and a prior encoding the $\ell_{1}$ constraint ibanez2023; Santucci2025; VAROQUAUX2013. Such an approach leads to a sparse, positive-definite precision matrix, which is interpretable as an adjacency matrix.